OpenStudy (anonymous):

Find the slope of the line segment containing the points (4, 3) and (-2, 7). i believe the answer is 3/2

OpenStudy (austinl):

\[m = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (austinl):

m = slope

OpenStudy (uri):

7-3/-2-4

OpenStudy (ivettef365):

you have 7 - 3 ----- -2 -4

terenzreignz (terenzreignz):

Well, there you have it, three responses, ranging from subtle, to strong, to direct. :D

OpenStudy (luigi0210):

Then after use the point-slope formula.

terenzreignz (terenzreignz):

^While that is good practice, I must point out that only the slope is required for this particular question :) Though, for future reference, the point-slope form is \[\Large y - y_0 = m(x-x_0)\]